OpenStudy (anonymous):

PLEASE help me with this simple problem. Me and my friend have been arguing for over an hour. I argue that -5^2 = 25. He argues that -5^2 = -25. Who is correct?

OpenStudy (anonymous):

Your friend.

OpenStudy (mathmate):

-5^2=-25 (-5)^2 = 25 It all boils down to the priority of operators.

OpenStudy (lgbasallote):

\[(-5)^2 = 25\] howeer.. \[-5^2 = -25\]

OpenStudy (lgbasallote):

lol beat me to it

OpenStudy (anonymous):

Wow.. Really? So -5^2 IS NOT -5*-5?

OpenStudy (mathmate):

Probably by a nanosecond!! :) But great minds think alike!

OpenStudy (lgbasallote):

nope

OpenStudy (lgbasallote):

-5^2 = -(5 * 5)

OpenStudy (anonymous):

I'm almost positive I've always learned differently, but thank you guys!

OpenStudy (lgbasallote):

the exponent is only for 5. the negative sign is not raised

OpenStudy (mathmate):

... unless parentheses are present.

OpenStudy (anonymous):

This needs to be closed